What is the capacitance of an empty parallel-plate capacitor with metal plates that each 2

have an area of 1.00 m , separated by 1.00 mm?. How much charge is stored in this

capacitor if a voltage of 3.00 x 103 V is applied to it?


1
Expert's answer
2022-03-13T18:46:06-0400

The capacitance of a parallel plate capacitor

"C=\frac{\epsilon_0A}{d}=\frac{8.85*10^{-12}*1.00}{1.00*10^{-3}}=8.85*10^{-9}\:\rm F"

The energy stored in capacitor

"W=\frac{CV^2}{2}\\
=\frac{8.85*10^{-9}\:\rm F*(3.00*10^3\:V)^2}{2}=0.0797\:\rm J"